# VideoPrism: A Foundational Visual Encoder for Video Understanding

### v1

@@ 5 Conclusion
Broader impacts
@@
Acknowledgements
We sincerely thank David Hendon for their product management efforts, and Alex Siegman, Ramya Ganeshan, and Victor Gomes for their program and resource management efforts. We also thank Hassan Akbari, Sherry Ben, Yoni Ben-Meshulam, Chun-Te Chu, Sam Clearwater, Yin Cui, Ilya Figotin, Anja Hauth, Sergey Ioffe, Xuhui Jia, Yeqing Li, Lu Jiang, Zu Kim, Dan Kondratyuk, Bill Mark, Arsha Nagrani, Caroline Pantofaru, Sushant Prakash, Cordelia Schmid, Bryan Seybold, Mojtaba Seyedhosseini, Amanda Sadler, Rif A. Saurous, Rachel Stigler, Paul Voigtlaender, Pingmei Xu, Chaochao Yan, Xuan Yang, and Yukun Zhu for the discussions, support, and feedback that greatly contributed to this paper. We thank Jay Yagnik, Rahul Sukthankar, and Tomas Izo for their enthusiastic support for this project.
@@ A.1 Data curation
Anonymous-Corpus #1.
This pretraining corpus consists of about 36M commercially licensed stock video-caption pairs, where the videos and text are manually uploaded by professional contributors. Hence, the quality of the videos and captions is high in this corpus compared with the rest.
Anonymous-Corpus #2.
This data corpus contains 170M (video, ASR transcript) pairs from 44.6M YouTube videos. Its construction process is similar to HowTo100M (Miech et al., 2019), but the whole corpus is larger and more diverse. Furthermore, the clip-text pairs are filtered based on a groundedness score similar to the CLIP’s similarity score (Wu et al., 2021).
Anonymous-Corpus #3.
This anonymous corpus includes 71.5M (clip, machine-generated caption) pairs from 36.7M YouTube videos. The clips are captioned using vision-language models (Chen et al., 2023c) and further summarized using an LLM (Anil et al., 2023). The corpus is similar to InternVid (Wang et al., 2023e) in terms of construction but a magnitude larger in size and diversity.
@@ Appendix D Evaluation data
Table 12: Results of FM adaptation using frozen features on video understanding tasks. The model backbones are frozen and only weights in the task heads are updated using the downstream tasks’ training sets. ${}^{\ast}$ indicates the model is evaluated under the setting with trainable FLOPs alignment.
@@
Table 13: Results of FM adaptation using frozen backbones with MLAP heads on video understanding tasks. MLAP takes multiple frozen features from an FM as inputs and map them hierarchically for the final task prediction. Only the MLAP layer weights are updated using the downstream tasks’ training sets. ${}^{\ast}$ indicates the model is evaluated under the setting with trainable FLOPs alignment.
@@
Table 14: Results of FM adaptation using frozen backbones with low-rank adapters and task heads. Only the weights of the low-rank adapters and task heads are updated using downstream tasks’ training sets. ${}^{\ast}$ indicates the model is evaluated under the setting with trainable FLOPs alignment.
@@
Table 15: Results of FM adaptation by end-to-end fine-tuning. All the model weights are updated using the downstream tasks’ training sets. ${}^{\ast}$ indicates the model is evaluated under the setting with trainable FLOPs alignment.
@@
Table 16: Stratified average scores under four adaptation methods and the final VGS. ${}^{\ast}$ indicates the model is evaluated under the setting with trainable FLOPs alignment.
@@ E.1 Tasks and task heads for VideoPrism
We experiment VideoPrism under two configurations regarding different input video sizes. In the first configuration (marked by asterisk “${}^{\ast}$” in Tables 12, 15, 14 and 13), we use $8$ frames and $252\times 252$ image resolution for feature extraction when training video classification task heads, which results in a sequence of $8\times 14\times 14$ tokens. On AVA and AVA-Kinetics, video clips of shape $8\times 288\times 288$ (i.e., token length $8\times 16\times 16$) are used for both training and evaluation. This configuration aligns the trainable FLOPs of VideoPrism with the other baseline models reported in Yuan et al. (2023). In the second configuration, we use video clips of shape $16\times 288\times 288$ as input for all the experiments. This configuration aligns with the pretraining setup with higher trainable FLOPs and accounts for the results in Section 3.1.
@@ E.3 Results
In Tables 12, 13, 14 and 15, we report the detailed benchmark results using the aforementioned four adaptation settings [^1] The UMT-B/16-25M checkpoint is obtained from https://github.com/OpenGVLab/unmasked_teacher/blob/main/multi_modality/MODEL_ZOO.md.. In Table 16, the stratified average scores for each FM under four adaptations are reported. We also report the final VideoGLUE score (VGS) for each FM, which weights their absolute performances with the respective adaptation costs.
Notably, from Tables 12, 13, 14 and 15, we can see that when evaluated under the same trainable FLOPs, VideoPrism consistently and significantly outperforms other FMs across different benchmarks and tasks. Aligning with the pretraining setup further improves the performance. The strong results under both configurations indicate the efficacy of the learned representations by VideoPrism.
@@ F.2 Zero-shot classification on Charades-STA
F.3 Additional results on Kinetics-600
@@ Appendix I Ablation studies
I.1 Data
We study how to combine datasets with different caption qualities, quantities, and distributions when training a video-text contrastive model. In Table 20, three different combination methods are considered: (1) simply mixing different datasets (denoted with “+” and “✗” for AGD); (2) training with one dataset and then continue training with another dataset (denoted with $\rightarrow$” and “✗” for AGD); (3) combining different methods with AGD (denoted with “+” and “✓” for AGD). We choose two representative datasets, namely InternVid (Wang et al., 2023e) and YTT180M (Zellers et al., 2021), and report Recall@1 (R@1) for zero-shot video-to-text (ZSV2T) and text-to-video (ZST2V) retrieval for this study. We notice that simply mixing InternVid and YTT180M results in a large performance drop on VATEX when compared with only using InternVid. On the other hand, training on one dataset then continue training on the other dataset is highly affected by the order of datasets. For instance, compared with only using InternVid, the performance of InternVid $\rightarrow$ YTT180M drops by a large margin on both MSRVTT and VATEX, while YTT180M $\rightarrow$ InternVid improves on three out of four metrics. Hence, this approach is not scalable with number of datasets. Alternatively, AGD consistently improves the performance on MSRVTT and ZSV2T of VATEX compared with YTT180M $\rightarrow$ InternVid and InternVid with only a slightly drop in ZST2V of VATEX. As a result, AGD is chosen to be used in combining different training datasets.
We further report the performance of AGD with all our pretraining corpus in the last row of Table 20. We observe a large improvement across all metrics with AGD, demonstrating that AGD scales well with the number of datasets.
@@ I.2 Model design
We first study the impact of masking method and masking ratio to the second-stage model using VideoPrism-B as an example. In Figure 6, we compare the performance of the second-stage model under tube masking (Tong et al., 2022) and BEVT masking (Wang et al., 2022b) with various masking ratios on different video focused tasks. We notice that BEVT masking outperforms tube masking in most cases. When comes to the masking ratio, BEVT masking with masking ratio $0.65$ and $0.75$ have similar performance and outperform the other masking ratios. As a result, if not otherwise specified, all the second-stage models are trained with BEVT masking with $0.65$ masking ratio.
Figure 7: Comparison between the first-stage and second-stage model of VideoPrism-B on video understanding tasks. For video-only tasks, results are from using MAP probing with frozen backbone.
@@
Figure 7 compares the performance of the second-stage model with the first-stage model on different video datasets across a variety of different tasks using VideoPrism-B. Specifically, for video only tasks such as video classification (K400 [VC] and SSv2 [VC]) and spatiotemporal action localization (AVA [STAL]), we apply the frozen feature and only update the weights in the task head. We report top-1 accuracy for video classification and mAP@IoU=$0.5$ for spatiotemporal action localization, respectively. For zero-shot video-to-text retrieval (MSRVTT [ZSV2T], VATEX [ZSV2T], and ActivityNet [ZSV2T]), zero-shot text-to-video retrieval (MSRVTT [ZST2V], VATEX [ZST2V], and ActivityNet [ZST2V]), and zero-short video classification (K400 [ZSC], K600 [ZSC], SSv2 [ZSC], and NExT-QA-ATP [ZSC]) tasks, we apply LiT (Zhai et al., 2022b) to learn a text encoder complied to the second-stage model as discussed in Section 3.2 and report Recall@1 (R@1). We notice that the second-stage training significantly improves the performance of the video encoder compared with the first-stage model across all video tasks on different datasets, strongly demonstrating the effectiveness of the proposed second-stage training.

### v3

@@ 5 Conclusion
Acknowledgements
We sincerely thank David Hendon for their product management efforts and the help with open-sourcing, and Alex Siegman, Ramya Ganeshan, and Victor Gomes for their program and resource management efforts. We also thank Hassan Akbari, Sherry Ben, Yoni Ben-Meshulam, Chun-Te Chu, Sam Clearwater, Yin Cui, Sathish Thoppay Egambaram , Ilya Figotin, Anja Hauth, Sergey Ioffe, Xuhui Jia, Lu Jiang, Zu Kim, Phoebe Kirk, Dan Kondratyuk, Yeqing Li, Bill Mark, Arsha Nagrani, Caroline Pantofaru, Sushant Prakash, Amanda Sadler, Rif A. Saurous, Cordelia Schmid, Bryan Seybold, Mojtaba Seyedhosseini, Rachel Stigler, Paul Voigtlaender, Pingmei Xu, Chaochao Yan, Xuan Yang, Rui Zhu, and Yukun Zhu for the discussions, support, and feedback that greatly contributed to this paper. Lastly, we thank Jay Yagnik, Rahul Sukthankar, and Tomas Izo for their enthusiastic support for this project.
Impact statement
@@ A.1 Data curation
Anonymous-Corpus #1 consists of about 36M commercially licensed stock video-caption pairs, where the videos and text are manually uploaded by professional contributors. Hence, the quality of the videos and captions is high in this corpus compared with the rest. Note that we do not do any filtering on this set.
Anonymous-Corpus #2 contains 170M (video, ASR transcript) pairs from 44.6M YouTube videos. Its construction process is similar to HowTo100M (Miech et al., 2019), but the whole corpus is larger and more diverse. Furthermore, view counts and video lengths are filtered using simple metadata. ASR sentence boundaries define the clip boundaries. The clip-text pairs are filtered based on a groundedness score similar to CLIP’s similarity score (Wu et al., 2021).
Anonymous-Corpus #3 includes 71.5M (clip, machine-generated caption) pairs from 36.7M YouTube videos. The clips are captioned using vision-language models (Chen et al., 2023c) and further summarized using an LLM (Anil et al., 2023). The corpus is similar to InternVid (Wang et al., 2023e) in terms of construction but a magnitude larger in size and diversity. The initial video selection of this dataset process is similar to Anonymous-Corpus #2, but additional filters are applied to exclude videos composed of primarily talking heads using a face detection model. Also, static clips are eliminated by ensuring semantic feature embeddings from the frames are not static. Hence, the video content is more diverse than Anonymous-Corpus #2.
@@ A.2 Corpus analysis
Furthermore, we provide an in-depth analyses on each Anonymous dataset and contrast them with the other datasets we used in Table 8. We find that the datasets with captions generated by VLMs and LLMs (e.g., InternVid and Anonymous-Corpus #3) perform the best. In addition, retrieval performance is correlated with dataset size, groundingness (CLIP score), “dynamic degree” (optical flow), and the presence of humans. Finally, we compute the top-50 object categories represented by our dataset obtained by running an open-source Tensorflow object detection API on the center-frame of 100K clips from our pretraining data. We find person to be the top category, followed by car, chair, TV, bottle, book, potted plant, and bowl. We hope this gives more insights into the composition of our datasets.
@@
Table 8: More details about different characteristics of each dataset. The noun/verb ratio and the presence of humans in the video are inferred from the captions using WordNet synsets (Fellbaum, 2005) and the NTLK library. Optical flow is computed at $8$ FPS and $256\times 256$ resolution, and the average optical flow across videos is presented. The last column, in addition to the data analyses, is the average of zero-shot R@1 text-video retrieval results on MSRVTT (full-split) and VATEX.
Datasets Caption source # of clips Noun / Verb % of videos Optical flow Grounding Average
ratio with humans magnitude (px) score ZS R@1
InternVid (Wang et al., 2023e) Generated 7.0M 3.4 79.9 2.47 0.32 45.7
YT-Temporal-180M (Zellers et al., 2021) ASR 87.8M 1.6 81.8 1.73 0.24 35.8
VideoCC (Nagrani et al., 2022) Retrievals 133.5M 8.0 69.0 3.21 0.27 32.6
WTS-70M (Stroud et al., 2020) Metadata 55.1M 1.3 18.5 3.67 0.26 19.4
Anonymous-Corpus #1 Manual 36.1M 5.9 76.9 1.55 0.30 37.0
Anonymous-Corpus #2 ASR 170.3M 2.8 62.2 2.15 0.26 38.6
Anonymous-Corpus #3 Generated 71.5M 3.2 92.0 3.17 0.30 49.4
@@ Appendix D Evaluation data
Table 13: Results of FM adaptation using frozen features on video understanding tasks. The model backbones are frozen and only weights in the task heads are updated using the downstream tasks’ training sets. indicates the model is evaluated under the setting with trainable FLOPs alignment.
@@
Table 14: Results of FM adaptation using frozen backbones with MLAP heads on video understanding tasks. MLAP takes multiple frozen features from an FM as inputs and map them hierarchically for the final task prediction. Only the MLAP layer weights are updated using the downstream tasks’ training sets. indicates the model is evaluated under the setting with trainable FLOPs alignment.
@@
Table 15: Results of FM adaptation using frozen backbones with low-rank adapters and task heads. Only the weights of the low-rank adapters and task heads are updated using downstream tasks’ training sets. indicates the model is evaluated under the setting with trainable FLOPs alignment.
@@
Table 16: Results of FM adaptation by end-to-end fine-tuning. All the model weights are updated using the downstream tasks’ training sets. indicates the model is evaluated under the setting with trainable FLOPs alignment.
@@
Table 17: Stratified average scores under four adaptation methods and the final VGS. indicates the model is evaluated under the setting with trainable FLOPs alignment.
@@ E.1 Tasks and task heads for VideoPrism
We experiment VideoPrism under two configurations regarding different input video sizes. In the first configuration (marked by asterisk “” in Tables 13, 16, 15 and 14), we use $8$ frames and $252\times 252$ image resolution for feature extraction when training video classification task heads, which results in a sequence of $8\times 14\times 14$ tokens. On AVA and AVA-Kinetics, video clips of shape $8\times 288\times 288$ (i.e., token length $8\times 16\times 16$) are used for both training and evaluation. This configuration aligns the trainable FLOPs of VideoPrism with the other baseline models reported in Yuan et al. (2023). In the second configuration, we use video clips of shape $16\times 288\times 288$ as input for all the experiments. This configuration aligns with the pretraining setup with higher trainable FLOPs and accounts for the results in Section 3.1.
@@ E.3 Results
In Tables 13, 14, 15 and 16, we report the detailed benchmark results using the aforementioned four adaptation settings [^1] The UMT-B/16-25M checkpoint is obtained from https://github.com/OpenGVLab/unmasked_teacher/blob/main/multi_modality/MODEL_ZOO.md.. In Table 17, the stratified average scores for each FM under four adaptations are reported. We also report the final VideoGLUE score (VGS) for each FM, which weights their absolute performances with the respective adaptation costs.
Notably, from Tables 13, 14, 15 and 16, we can see that when evaluated under the same trainable FLOPs, VideoPrism consistently and significantly outperforms other FMs across different benchmarks and tasks. Aligning with the pretraining setup further improves the performance. The strong results under both configurations indicate the efficacy of the learned representations by VideoPrism.
@@
Table 18: Zero-shot video-text retrieval on MSRVTT. We follow the full split produced by Xu et al. (2016) which contains $2,990$ videos for testing.
Methods MSRVTT
Text $\rightarrow$ Video Video $\rightarrow$ Text
R@1 R@5 R@1 R@5
CLIP-B (Radford et al., 2021) 23.3 44.2 43.3 73.3
SM-B (Zeng et al., 2022) - - 46.9 73.5
CoCa-g (Yu et al., 2022) 30.0 52.4 49.9 73.4
VideoCoCa-g (Yan et al., 2022) 34.3 57.8 64.7 85.2
VideoPrism-B 37.0 61.5 67.7 87.5
($\uparrow$2.7) ($\uparrow$3.7) ($\uparrow$3.0) ($\uparrow$2.3)
VideoPrism-g 39.7 63.7 71.0 90.0
($\uparrow$5.4) ($\uparrow$5.9) ($\uparrow$6.3) ($\uparrow$4.8)
@@ F.2 Zero-shot classification on Charades-STA
F.3 Additional results on MSRVTT
In Table 18, we report zero-shot video-text retrieval results on the full split of MSRVTT produced by Xu et al. (2016), which contains $2,990$ videos for testing. We observe that VideoPrism outperforms previous methods by a large margin. More importantly, our base-scale model is also better than existing larger-scale models (e.g., CoCa-g and VideoCoCa-g). These findings are consistent with the results in Table 3, which confirms the strong capability of VideoPrism on zero-shot video-text retrieval tasks. Note that, in the appendix, all video-text retrieval results on MSRVTT are calculated using this full split, unless otherwise stated.
F.4 Additional results on Kinetics-600
@@ Appendix I Ablation studies
I.1 Data
We study how to combine datasets with different caption qualities, quantities, and distributions when training a video-text contrastive model. In Table 22, three different combination methods are considered: (1) simply mixing different datasets (denoted with “+” and “✗” for AGD); (2) training with one dataset and then continue training with another dataset (denoted with $\rightarrow$” and “✗” for AGD); (3) combining different datasets with AGD (denoted with “+” and “✓” for AGD). We choose two representative datasets, namely InternVid (Wang et al., 2023e) and YTT180M (Zellers et al., 2021) and report Recall@1 (R@1) for zero-shot video-to-text (ZSV2T) and text-to-video (ZST2V) retrieval for this study. We notice that simply mixing InternVid and YTT180M results in a large performance drop on VATEX when compared with only using InternVid. On the other hand, training on one dataset then continue training on the other dataset is highly affected by the order of datasets. For instance, compared with only using InternVid, the performance of InternVid $\rightarrow$ YTT180M drops by a large margin on both MSRVTT and VATEX, while YTT180M $\rightarrow$ InternVid improves on three out of four metrics. Hence, this approach is not scalable with the number of datasets. Alternatively, AGD consistently improves the performance on MSRVTT and ZSV2T of VATEX compared with YTT180M $\rightarrow$ InternVid and InternVid with only a slightly drop in ZST2V of VATEX. As a result, AGD is chosen for combining different training datasets when we train the video-text contrastive model.
We further report the performance of AGD with all our pretraining corpus in the last row of Table 22. We observe a large improvement across all metrics, demonstrating that AGD scales well with the number of datasets.
@@ I.1 Data
To better understand how our models perform when only public datasets are included for pretraining, we conduct experiments under two setups where only (1) InternVid (7M) and (2) all public datasets (150M, including InternVid, YT-Temporal-180M, and WTS-70M) are used, respectively. As shown in Table 23, we find our first-stage model has already achieved overall favorable results, while our second-stage model yields better results, especially on SSv2. As with the case with any foundation model, the pretraining data is one of the factors to improve performance, but it is not the only factor. The proposed two-stage pretraining strategy, token-wise distillation with token shuffling, and global distillation all contribute to the performance improvements.
Table 23: Model performance when only public datasets are used for pretraining. Results using MAP probing with frozen backbone for VideoPrism-B are reported. We report Top-1 accuracy on K400 and SSv2.
Methods Corpus size K400 SSv2
UMT-B (Li et al., 2023b) 25M 77.1 47.7
VideoMAE-B (Tong et al., 2022) 1M 65.1 53.9
VideoPrism-B (Stage 1) 7M 81.0 47.6
VideoPrism-B (Stage 1) 150M 81.7 50.0
VideoPrism-B (Stage 2) 7M 81.5 60.3
VideoPrism-B (Stage 2) 150M 82.7 60.5
@@ I.2 Model design
Table 24: Joint vs. factorized attention. Results of the first-stage VideoPrism-B model are reported. We report Top-1 accuracy on K400 and SSv2 using MAP probing with frozen backbone.
Encoder architectures MSRVTT K400 SSv2
ZST2V ZSV2T VC VC
Joint attention 34.2 64.0 84.5 52.6
Factorized attention 34.2 64.6 82.9 55.5
Factorized encoder.
We favored the factorized attention over joint attention in the encoder because it balances the cost (e.g., memory, efficiency) and performance well. The factorized attention is especially appealing given that contrastive learning (our Stage 1) demands a large batch size. We note that the original ViViT paper (Arnab et al., 2021) also recommended the factorized-attention architecture over the other variants. In Table 24, we can see that the two attention schemes lead to similar performance on text-video retrieval tasks, while the joint attention gives rise to slightly better accuracy on K400 and yet worse on SSv2. The overall performances of these two models are similar, reinforcing that the factorized attention is probably a better choice given its smaller memory footprint.
Table 25: Effects of model initialization. Results of the first-stage VideoPrism-B model are reported. We report Top-1 accuracy on K400 and SSv2 using MAP probing with frozen backbone.
Configurations MSRVTT K400 SSv2
ZST2V ZSV2T VC VC
CLIP (Radford et al., 2021) (frozen) 27.6 52.4 73.1 39.8
CLIP (Radford et al., 2021) 27.4 53.3 79.6 51.1
CoCa (Yu et al., 2022) 29.9 57.3 81.7 50.0
Initialization.
To understand how the end results vary with respect to the initialization from different image models, we replace CoCa (Yu et al., 2022) with CLIP (Radford et al., 2021) as the spatial encoder in our model. We then experiment with two variants: (1) freezing CLIP while training the temporal layers and (2) unfreezing all weights in training. Note that we use CLIP-B/14 and all the public datasets in our pretraining corpus for model training in this experiment. Results are shown in Table 25. We can see that the results using CLIP for initialization are still comparable with those with the CoCa initialization, and unfreezing the CLIP weights benefits all evaluation benchmarks in the table.
@@
We then study the impact of masking method and masking ratio on the second-stage model using VideoPrism-B as an example. In Figure 6, we compare the performance of the second-stage model under tube masking (Tong et al., 2022) and BEVT masking (Wang et al., 2022b) with various masking ratios on different video focused tasks. We notice that BEVT masking outperforms tube masking in most cases. When comes to the masking ratio, BEVT masking with masking ratio $0.65$ and $0.75$ have similar performance and outperform the other masking ratios. As a result, if not otherwise specified, all the second-stage models are trained with the BEVT masking with $0.65$ masking ratio.
Figure 7: Comparison between the first-stage and second-stage models of VideoPrism-B on video understanding tasks. For video-only tasks, results are from using MAP probing with a frozen backbone.
@@
Figure 7 compares the second-stage model with the first-stage model on different video datasets across a variety of tasks using VideoPrism-B. Specifically, for video-only tasks such as video classification (K400 [VC] and SSv2 [VC]) and spatiotemporal action localization (AVA [STAL]), we apply the frozen feature and only update the weights in the task head. We report top-1 accuracy for video classification and mAP@IoU=$0.5$ for spatiotemporal action localization, respectively. For zero-shot video-to-text retrieval (MSRVTT [ZSV2T], VATEX [ZSV2T], and ActivityNet [ZSV2T]), zero-shot text-to-video retrieval (MSRVTT [ZST2V], VATEX [ZST2V], and ActivityNet [ZST2V]), and zero-short video classification (K400 [ZSC], K600 [ZSC], SSv2 [ZSC], and NExT-QA-ATP [ZSC]) tasks, we apply LiT (Zhai et al., 2022b) to learn a text encoder paired up with the second-stage model as described in Section 3.2 and report Recall@1 (R@1). We notice that the second-stage training significantly improves the performance of the video encoder compared with the first-stage model across all video tasks on different datasets, strongly demonstrating the effectiveness of the proposed two-stage training.